1. The visible light spectrum has a wavelength range of 380-700nm. Find the frequency of a wave with the wavelength of 410nm. (Hint: the speed of light is 3x108 m/s) 

 


1
Expert's answer
2021-08-07T12:37:42-0400

Given:

"\lambda=410\times 10^{-9}\:\rm m"

"c=3\times 10^{8}\:\rm m/s"

The frequency

"f=\frac{c}{\lambda}=\frac{3\times 10^8}{410\times 10^{-9}}=7.32\times 10^{14}\:\rm Hz"